Stock X has a 9.5% expected return, a beta coefficient of 0.8, and a 35% standard deviation of expected returns. Stock Y hasc. Calculate each stocks required rate of return. Round your answers to two decimal places. rx = % ry %e. Calculate the required return of a portfolio that has $4,000 invested in Stock X and $8,000 invested in Stock Y. Do not roI will rate! Thanks in advance!
Answer #1

a). Calculating the Stock's Coefficient of variation:-

Coefficient of variation = Standard Deviation/Expected return

CV_x = \frac{35}{9.5}

= 3.68

CV_y = \frac{25}{12}

= 2.08

c). As per CAPM,

Required rate of Return = R-f + \beta (R_{mp})

Rf = Risk Free Return =6%

Rmp = Market Risk Premium = 5%

- Beta of Stock A = 0.8

Required rate of retrun = 6% +0.8(5%)

= 10%

- Beta of Stock B = 1.1

Required rate of retrun = 6% +1.1(5%)

= 11.5%

e). You Invested $4000 in Stock X

& Invested $ 8000 in Stock Y

Calculating the Required return of Portfolio:-

Required return of Portfolio = (Weight of X in Portfolio)(Required Return of X) + (Weight of Y in Portfolio)(Required Return of Y)

= [4000/(4000+8000)]*(10%) + [8000/(4000+8000)]*(11.5%)

= 3.33% + 7.67%

= 11%